# The Collection of a Human Robot Collaboration Dataset for Cooperative Assembly in Glovebox Environments

### v1

@@ A.2 Overview of Hand Segmentation Datasets
GTEA [15]  31K 663 28 34 m 4 1280 x 720 act msk
ADL [27] >1M 1,000,000 20  10 h 20 1280 x 720 act obj
EDSH [22]  20K 442 3  10 m - 1280 x 720 msk
Interactive Museum [2] - 700 - 5 800 x 450 gst msk
EgoHands [1]  130K 4800 48 72 m 8 1280 x 720 msk
Maramotti [3] - - 700 - 5 800 x 450 gst msk
UNIGE Hands [4]  150K - - 98 m - 1280 x 720 det
GUN-71 [29]  12K 9100 - - 8 - grs
RGBD Egocentric Action [35] - - - - 20 C: 640x480 D: 320 x 240 act
Fingerwriting in mid-air [8]  8K 2500 - - - - ftp gst
Ego-Finger [19]  93K - 24 - - 640 x 480 det ftp
ANS able-bodied [25] - - - 44 m 4 640 x 480 det ftp
UT Grasp [6] - - 50  4 h 5 960 x 540 grs
GestureAR [26]  51K - 100 - 8 1280 x 720 gst
EgoGesture [37]  59K - - - - - det ftp gst
Egocentric hand-action [38]  154K - 300 - 26 320 x 240 gst
BigHand2.2M [39]  290K  290K - - - 640 x 480 pos
Desktop Action [7]  324K 660 60 3 h 6 1920 x 1080 act msk
Epic Kitchens [11] 11.5M - - 55h 32 1920 x 1080 act
FPHA [16]  100K  100K 1175 - 6 C: 1920 x 1080 D: 640 x 480 act pos
EYTH [21] 1290 1290 3 - - 384 x 216 msk
HandOverFace [21] 300 300 - - - 384 x 216 msk
EGTEA+ [23] >3M  15k 86  28h 32 1280 x 960 act gaz msk
THU-READ [33]  343K 652 1920 - 50 640 x 480 act msk
EgoGesture [41]  3M - 24161 - 50 640 x 480 gst
HOI [20] 27525 - - - 6 - msk
EgoDaily [10]  50K  50K 50 - 10 1920 x 1080 det hid
ANS SCI [24] -  33K - - 17 480 x 854 det int
KBH [36]  12.5K  12.5K 161 - 50 230 x 306 msk
WorkingHands [32] 4.2K syn, 3.7K Real  7.8k 39 - - 1920 × 1080 obj msk
Freihand [42] 37000 37000 - - 32 - pos
MECCANO [28]  299K - 20 21 min 20 1280 x 720 act obj msk
HRC [30] 598 - 13 - 2 380 x 180 (resized) msk
HaDR [18] 117000 117000 - - - 640 x 480 msk
@@
GTEA [15] Pixel N 17 16
ADL [27] Bounding Box N 11 10
EDSH [22] Pixel N 1 0
Interactive Museum [2] Bounding Box N 1 0
EgoHands [1] Pixel N 4 0
Maramotti [3] Pixel N 1 0
UNIGE Hands [4] - N 1 0
GUN-71 [29] Key Points + Forces N 1 0
RGBD Egocentric Action [35] - N 1 0
Fingerwriting in mid-air [8] Key Point N 1 0
Ego-Finger [19] Bounding Box + Points N 1 0
ANS able-bodied [25] Bounding Box N 1 0
UT Grasp [6] Bounding Box N 1 0
GestureAR [26] Pixel 1 0
EgoGesture [37] Bounding Box + Points N 1 0
Egocentric hand-action [38] Bounding Box Y 1 0
BigHand2.2M [39] Key Point N 1 0
Desktop Action [7] Pixel N 1 0
Epic Kitchens [11] Bounding Box N 324 323
FPHA [16] Key Point N 1 0
EYTH [21] Pixel N 1 0
HandOverFace [21] Pixel N 1 0
EGTEA+ [23] Pixel N 1 0
THU-READ [33] Pixel N 1 0
EgoGesture [41] Pixel N 1 0
HOI [20] Bounding Box N 1 0
EgoDaily [10] Bounding Box N 1 0
ANS SCI [24] Pixel N 1 0
KBH [36] Pixel N 1 0
WorkingHands [32] Pixel Y 14 13
Freihand [42] Key Point N 1 0
MECCANO [28] Bounding Box N 21 20
HRC [30] Pixel N 5 1
HaDR [18] Pixel Y 1 0

### v2

@@ 5.2 Overview of Hand Segmentation Datasets
GTEA (Fathi et al., 2011)  31K 663 28 34 m 4 1280 x 720 act msk
ADL (Pirsiavash and Ramanan, 2012) ¿1M 1,000,000 20  10 h 20 1280 x 720 act obj
EDSH (Li and Kitani, 2013)  20K 442 3  10 m - 1280 x 720 msk
Interactive Museum (Baraldi et al., 2014) - 700 - 5 800 x 450 gst msk
EgoHands (Bambach et al., 2015)  130K 4800 48 72 m 8 1280 x 720 msk
Maramotti (Baraldi et al., 2015) - - 700 - 5 800 x 450 gst msk
UNIGE Hands (Betancourt et al., 2015)  150K - - 98 m - 1280 x 720 det
GUN-71 (Rogez et al., 2015)  12K 9100 - - 8 - grs
RGBD Egocentric Action (Wan and Aggarwal, 2015) - - - - 20 C: 640x480 D: 320 x 240 act
Fingerwriting in mid-air (Chang et al., 2016)  8K 2500 - - - - ftp gst
Ego-Finger (Huang et al., 2016)  93K - 24 - - 640 x 480 det ftp
ANS able-bodied (Likitlersuang and Zariffa, 2018) - - - 44 m 4 640 x 480 det ftp
UT Grasp (Cai et al., 2017) - - 50  4 h 5 960 x 540 grs
GestureAR (Mohatta et al., 2017)  51K - 100 - 8 1280 x 720 gst
EgoGesture (Wu et al., 2017)  59K - - - - - det ftp gst
Egocentric hand-action (Xu et al., 2016)  154K - 300 - 26 320 x 240 gst
BigHand2.2M (Yuan et al., 2017)  290K  290K - - - 640 x 480 pos
Desktop Action (Cai et al., 2019)  324K 660 60 3 h 6 1920 x 1080 act msk
Epic Kitchens (Damen et al., 2018) 11.5M - - 55h 32 1920 x 1080 act
FPHA (Garcia-Hernando et al., 2018)  100K  100K 1175 - 6 C: 1920 x 1080 D: 640 x 480 act pos
EYTH (Khan and Borji, 2018) 1290 1290 3 - - 384 x 216 msk
HandOverFace (Khan and Borji, 2018) 300 300 - - - 384 x 216 msk
EGTEA+ (Li et al., 2020) ¿3M  15k 86  28h 32 1280 x 960 act gaz msk
THU-READ (Tang et al., 2019)  343K 652 1920 - 50 640 x 480 act msk
EgoGesture (Zhang et al., 2018)  3M - 24161 - 50 640 x 480 gst
HOI (Kang et al., 2018) 27525 - - - 6 - msk
EgoDaily (Cruz and Chan, 2019)  50K  50K 50 - 10 1920 x 1080 det hid
ANS SCI (Likitlersuang et al., 2019) -  33K - - 17 480 x 854 det int
KBH (Wang et al., 2019)  12.5K  12.5K 161 - 50 230 x 306 msk
WorkingHands (Shilkrot et al., 2019) 4.2K syn, 3.7K Real  7.8k 39 - - 1920 × 1080 obj msk
Freihand (Zimmermann et al., 2019) 37000 37000 - - 32 - pos
MECCANO (Ragusa et al., 2020)  299K - 20 21 min 20 1280 x 720 act obj msk
HRC (Sajedi et al., 2022) 598 - 13 - 2 380 x 180 (resized) msk
HaDR (Grushko et al., 2023) 117000 117000 - - - 640 x 480 msk
@@
GTEA (Fathi et al., 2011) Pixel N 17 16
ADL (Pirsiavash and Ramanan, 2012) Bounding Box N 11 10
EDSH (Li and Kitani, 2013) Pixel N 1 0
Interactive Museum (Baraldi et al., 2014) Bounding Box N 1 0
EgoHands (Bambach et al., 2015) Pixel N 4 0
Maramotti (Baraldi et al., 2015) Pixel N 1 0
UNIGE Hands (Betancourt et al., 2015) - N 1 0
GUN-71 (Rogez et al., 2015) Key Points + Forces N 1 0
RGBD Egocentric Action (Wan and Aggarwal, 2015) - N 1 0
Fingerwriting in mid-air (Chang et al., 2016) Key Point N 1 0
Ego-Finger (Huang et al., 2016) Bounding Box + Points N 1 0
ANS able-bodied (Likitlersuang and Zariffa, 2018) Bounding Box N 1 0
UT Grasp (Cai et al., 2017) Bounding Box N 1 0
GestureAR (Mohatta et al., 2017) Pixel 1 0
EgoGesture (Wu et al., 2017) Bounding Box + Points N 1 0
Egocentric hand-action (Xu et al., 2016) Bounding Box Y 1 0
BigHand2.2M (Yuan et al., 2017) Key Point N 1 0
Desktop Action (Cai et al., 2019) Pixel N 1 0
Epic Kitchens (Damen et al., 2018) Bounding Box N 324 323
FPHA (Garcia-Hernando et al., 2018) Key Point N 1 0
EYTH (Khan and Borji, 2018) Pixel N 1 0
HandOverFace (Khan and Borji, 2018) Pixel N 1 0
EGTEA+ (Li et al., 2020) Pixel N 1 0
THU-READ (Tang et al., 2019) Pixel N 1 0
EgoGesture (Zhang et al., 2018) Pixel N 1 0
HOI (Kang et al., 2018) Bounding Box N 1 0
EgoDaily (Cruz and Chan, 2019) Bounding Box N 1 0
ANS SCI (Likitlersuang et al., 2019) Pixel N 1 0
KBH (Wang et al., 2019) Pixel N 1 0
WorkingHands (Shilkrot et al., 2019) Pixel Y 14 13
Freihand (Zimmermann et al., 2019) Key Point N 1 0
MECCANO (Ragusa et al., 2020) Bounding Box N 21 20
HRC (Sajedi et al., 2022) Pixel N 5 1
HaDR (Grushko et al., 2023) Pixel Y 1 0